3) (5 points) In the circuit below: a) Determine all the currents b) Find Vab 18.0 V 8.00 Ω 12.0 V 36.0 V
Answer #1

writing loop equation for the top loop we have

18 - 8I1 - 11i2 + 12 = 0

8I1 + 11I2 = 30 ---------(1)

writing loop equation for bottom wire we have

-11I2 + 12 - 36 = 0

I2 = -2.18 A

8I1 + 11*-2.18 = 30

I1 = 6.75 A

from the figure

I1 = I2 + I3

I3 = I1 - I2 = 6.75 - (-2.18) = 8.93 A

b) Va + 18 - 8*6.75 = Vb

Va - Vb = 36 V